CALL FOR PAPERS

LOPSTR’98

8th International Workshop on Logic-based Program Synthesis and Transformation

Manchester, United Kingdom, 15 – 19 June 1998

Sponsored by the Association for Logic Programming (ALP) and by the ESPRIT Network of Excellence in Computational Logic (Compulog Net)

LOPSTR’98 is the eighth in a series of international workshops on logic-based program synthesis and transformation. It will be held in parallel to the Joint International Conference and Symposium on Logic Programming (JICSLP’98).

New Objectives

As a workshop, LOPSTR’98 will continue the tradition of being a lively and friendly forum for presenting recent and current research as well as discussing future trends in the synthesis and transformation of programs. Compared to previous editions, however, the scope and mode of operation are somewhat modified. The scope now encompasses any computational-logic-based techniques, languages, and tools for the interactive or automated development of any kinds of programs. Also, papers discussing programming-in-the-large issues, or presenting practical applications, or convincingly arguing for the practical applicability of given theoretical results are now strongly encouraged. The mode of operation now is as follows. Based upon submitted extended abstracts, the programme committee will invite authors to present their research at the workshop. Pre-proceedings with the accepted extended abstracts will be available at the workshop as a technical report. Shortly after the workshop, the programme committee will invite the authors of the most promising abstracts and presentations to submit full papers. After another round of refereeing, the best full papers will be included in the post-workshop proceedings, which are to be published within the Lecture Notes in Computer Science series of Springer-Verlag. At this point, extended abstracts of at most eight pages are thus solicited about, but not limited to, the following topics:

specification  composition  specialization  verification  schemas
synthesis  transformation  analysis  reuse  industrial applications

Every submission must clearly exhibit the relationship to the (new) scope of the workshop, and must really be an extended abstract. It must thus be a miniature research paper with the key motivations and ideas, with outlines of the proofs of the key theorems, with references and a (favorable) comparison to related work, but without full details of proofs or implemented systems, without the description of future work, without ramifications that are irrelevant to the key ideas. A good extended abstract convinces a specialist referee of the significance of the described research in about one hour. Submissions that fail to comply with these content and size requirements may be rejected (without refereeing), no matter how high the scientific merit of the discussed work.

Programme Committee

Pierre Flener, Bilkent University, Turkey (programme chair)

Nicoletta Cocco, University of Venice, Italy  Ali Mili, Institute for Software Research, USA
Andreas Hamfelt, Uppsala University, Sweden  Lee Naish, Melbourne University, Australia
Kung-Kiu Lau, University of Manchester, UK (workshop chair)  Mario Ornaghi, University of Milan, Italy
Baudouin Le Charlier, University of Namur, Belgium  Alberto Pettorossi, University of Rome II, Italy
Michael Leuschel, Catholic University of Leuven, Belgium  Dave Robertson, University of Edinburgh, UK
Michael Lowry, NASA Ames, USA  Richard Waldinger, SRI International, USA

Important Dates

Submission of extended abstracts: 27 March 1998  Submission of full papers: 25 September 1998
Notification to authors:  1 May 1998  Notification to authors:  30 October 1998

Submissions

Electronic submissions are strongly encouraged. See the workshop homepage for full instructions regarding the formatting requirements and the electronic submission procedure. Address all correspondence (email preferred) to:

Pierre Flener (lopstr’98)  Phone:  +90 312 266 4000 ext. 1450 (gmt+2)
Department of Computer Science  Fax:  +90 312 266 4126
Bilkent University  Email:  {pf,lopstr98}@cs.bilkent.edu.tr
06533 Ankara  Homepage: http://www.cs.bilkent.edu.tr/∼lopstr98
Turkey  EU mirror: http://www.cs.man.ac.uk/∼kung-kiu/lopstr98